What is the approximate mass of Ca(NO3​)2​?(Given- Atomic mass of Ca=40,N=14,O=16)

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Jayesh Vashishtha
2 years ago
Dear student,
 
The approximate mass of Ca(NO3​)2​ =40+2×[14+(3×16)]=40+2×(14+48)=40+2×62=40+124=164. 
 
regards,
Team Askiitians